### v1

# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

Abstract
Histopathology and transcriptomics are fundamental modalities in cancer diagnostics, encapsulating the morphological and molecular characteristics of the disease. Multi-modal self-supervised learning has demonstrated remarkable potential in learning pathological representations by integrating diverse data sources. Conventional multi-modal integration methods primarily emphasize modality alignment, while paying insufficient attention to retaining the modality-specific intrinsic structures. However, unlike conventional scenarios where multi-modal inputs often share highly overlapping features, histopathology and transcriptomics exhibit pronounced heterogeneity, offering orthogonal yet complementary insights. Histopathology data provides morphological and spatial context, elucidating tissue architecture and cellular topology, whereas transcriptomics data delineates molecular signatures through quantifying gene expression patterns. This inherent disparity introduces a major challenge in aligning these modalities while maintaining modality-specific fidelity. To address these challenges, we present MIRROR, a novel multi-modal representation learning framework designed to foster both modality alignment and retention. MIRROR employs dedicated encoders to extract comprehensive feature representations for each modality, which is further complemented by a modality alignment module to achieve seamless integration between phenotype patterns and molecular profiles. Furthermore, a modality retention module safeguards unique attributes from each modality, while a style clustering module mitigates redundancy and enhances disease-relevant information by modeling and aligning consistent pathological signatures within a clustering space. Extensive evaluations on The Cancer Genome Atlas (TCGA) cohorts for cancer subtyping and survival analysis highlight MIRROR’s superior performance, demonstrating its effectiveness in constructing comprehensive oncological feature representations and benefiting the cancer diagnosis.
Index Terms:
Pathology, Whole Slide Image (WSI), Transcriptomics, Self-Supervised Learning (SSL), Multimodal Learning
I Introduction
Figure 1: MIRROR compared with conventional multi-modal integration methods. Unlike conventional methods that primarily emphasize capturing modality-shared information while paying limited attention to modality-specific intrinsic structures and indiscriminately learning both disease-relevant and irrelevant data with high redundancy, MIRROR is specifically designed to balance modality alignment and retention. By selectively preserving only disease-relevant features, it effectively mitigates redundancy, thereby enhancing the model’s efficiency and representational capability.
Histopathology images are widely regarded as the gold standard in cancer diagnosis, offering critical insights into the presence, type, grade, and prognosis of cancer [1]. These images encapsulate a wealth of morphological features that serve as the foundation of cancer diagnostics [2, 3, 4, 5]. Meanwhile, advancements in high-throughput sequencing technologies, such as polymerase chain reaction (PCR) [6], have further expanded oncological diagnostic capabilities by enabling the analysis of molecular data, including transcriptomics data, which delineates gene expression profiles, offering molecular signatures of the disease. The integration of the morphological and molecular modalities within multi-modal diagnostics significantly enhances the accuracy of cancer diagnosis and prognosis prediction, providing a more comprehensive and precise understanding of the disease.
Despite the transformative potential of multi-modal diagnostics, several challenges hinder its widespread adoption. The incorporation of molecular data into existing pathology workflows presents substantial complexities, further exacerbating the workload for already overburdened pathologists. Additionally, the scarcity of annotated paired data, due to the resource-intensive and time-consuming nature of labeling, significantly hinders the adoption of supervised learning methods. In this context, multi-modal self-supervised learning (SSL) emerges as a promising alternative, offering the capability to capture robust and comprehensive oncological feature representations without reliance on extensive annotations. Multi-modal SSL methods [7, 8, 9] have demonstrated remarkable success in both natural and medical domains, effectively aligning modalities such as image-text pairs. However, the direct extension of such aligning techniques to histopathology and transcriptomics data presents distinct challenges. Unlike conventional use cases where multi-modal inputs often exhibit highly overlapping features, histopathology and transcriptomics data pairs are inherently more heterogeneous, as they operate at different biological scales and encode distinct yet complementary dimensions of disease-related information. Histopathology provides a morphological and spatial view of tissue architecture, capturing phenotypic traits, while transcriptomics quantifies gene expression levels, uncovering the molecular processes and pathways underlying the disease. Although there are shared correlations between these modalities, each modality also retains substantial modality-specific information [1, 10]. As shown in Figure 1, existing multi-modal methods [11, 12, 13, 14] primarily focus on aligning shared information between modalities, while giving comparatively less attention to the rich and modality-specific information inherent to each data type. For instance, in [14], the authors employed an encoder-only architecture with contrastive learning to enforce representation alignment. However, the optimum training target will be achieved when representations from the same sample become indistinguishable across modalities, thereby eliminating essential modality-specific attributes. Moreover, both histopathology and transcriptomics data contain redundant, disease-unrelated information, including repetitive structural patterns and genes with overlapping functions or pathways. Reducing such redundancies can enable the model to extract more clinically meaningful, disease-relevant representations. Beyond these challenges, the inherent heterogeneity in data formats presents an additional layer of complexity. The histopathology data is structured as 2D image patches, whereas transcriptomic data is represented as tabular numerical values, necessitating careful architectural design.
To address these challenges, we propose MIRROR (Multi-modal pathologIcal self-supeRvised Representation learning via mOdality alignment and Retention), a novel multi-modal SSL framework designed to foster both modality alignment and retention. MIRROR adopts dedicated Transformer-based [15] encoders to extract rich and discriminative feature representations for each modality while being specifically tailored to accommodate the inherent heterogeneity in data format. To facilitate seamless integration within the latent space, a modality alignment module is introduced, dynamically drawing paired data into closer proximity while dispersing unrelated samples. To safeguard modality-specific fidelity, a modality retention module is employed to ensure the preservation of unique modality attributes. This module challenges the model to maintain modality-specific intrinsic structures by reconstructing key features after perturbation. Additionally, to mitigate redundancy and enhance disease-relevant information, MIRROR incorporates a style clustering module, which maps feature embeddings into a statistical space to capture consistent pathological styles while minimizing intra-modality redundancy. Subsequently, a prototype clustering mechanism further aligns the captured styles in the clustering space, mitigating inter-modality redundancy and reinforcing biologically meaningful correspondences. Together, these synergistic modules cultivate a well-structured representation space by disentangling and preserving both modality-shared and modality-specific signatures while suppressing irrelevant variations, enabling MIRROR to deliver robust and comprehensive multi-modal representations, advancing the capabilities of multi-modal diagnostics.
Furthermore, the vast number of genes available in transcriptomics data presents a significant challenge in identifying those most relevant to disease development. MIRROR addresses this issue through a novel preprocessing pipeline that integrates both machine learning-driven feature selection with biological knowledge to distill high-dimensional transcriptomics data, creating refined and disease-focused transcriptomics datasets.
The proposed method is evaluated using 5-fold cross-validation on multiple cohorts from the TCGA dataset [16], targeting critical downstream tasks including cancer subtyping and survival analysis. The evaluation incorporates both linear probing and few-shot learning settings to comprehensively assess the performance and generalizability of the model. The key contributions of this study are outlined as follows:
  • MIRROR, a novel multi-modal SSL model, is designed to facilitate both modality alignment and retention, enabling the effective preservation of both modality-shared and modality-specific information.
  • A consistent pathological style-based clustering mechanism is introduced to preserve disease-relevant information while mitigating redundancy.
  • A novel preprocessing pipeline for transcriptomics data is proposed, integrating machine learning-driven feature selection with biological knowledge to create refined transcriptomics datasets.
  • Comprehensive evaluations are conducted across diverse cohorts from the TCGA dataset, focusing on cancer subtyping and survival analysis tasks, substantiating the superior performance and effectiveness of the proposed approach.
II Related Work
II-A Vision Self-Supervised Learning
Recent advancements in SSL in computer vision (CV) have significantly reduced the reliance on labeled data while achieving performance comparable to, or even surpassing, that of supervised methods. SSL can be categorized into three main styles: contrastive learning, generative learning, and hybrid methods that integrate both approaches. Contrastive learning models [17, 18] achieve representation learning by drawing closer the embeddings of the same image under different augmentations while separating those of different images. Generative approaches [19, 20] aim to reconstruct or predict missing or unobserved portions of the input using the information from the observed parts. Hybrid methods, such as [21, 22], integrate the strengths of contrastive and generative learning to simultaneously capture both discriminative and semantic information effectively.
Multi-modal SSL has further enhanced the capability of models to integrate and align data from different modalities. Among various inputs, vision-language learning [7, 23, 24] is the most extensively studied and demonstrates exceptional performance. Due to the largely overlapping information between image and text data, aligning the two modalities in the latent space alone is often sufficient to achieve impressive results.
II-B Pathological Self-Supervised Learning
To overcome the scarcity of labeled data in Computational Pathology (CPath), a large amount of prior works [25, 26, 27, 28, 29] has explored unsupervised learning and transfer learning techniques. In recent years, SSL methods have gained increasing recognition for their superior potential in CPath, particularly in leveraging large-scale unlabeled Whole Slide Images (WSIs). Similar to general CV, SSL approaches in CPath can be broadly classified into three categories: contrastive [3, 30, 31, 32, 33], generative [34, 35], and hybrid [4, 36, 37, 38].
Multi-modal pathological SSL enhances the model’s ability to learn comprehensive pathological representations by integrating diverse inputs from the diagnostic process. This includes combining WSIs with clinical reports or text captions [39, 40, 41], incorporating transcriptomics data [11], and fusing images from different staining techniques [35, 42]. However, these methods often overlook the inherent redundancy in WSIs and other inputs, leading to the indiscriminate learning of both disease-relevant and disease-irrelevant information, which hampers the models’ performance and efficiency.
II-C Histopathology and Transcriptomics Multi-Modal Learning
Multi-modal integration methods [12, 43, 44, 11, 45, 14, 46] that combine histopathology and transcriptomics have shown impressive capabilities in various pathological tasks, including cancer diagnosis and survival analysis. While these modalities share some information, they also contain a vast amount of comprehensive and distinct insights into the disease, reflecting their inherent heterogeneity. Despite improvements in performance, existing multi-modal methods predominantly focus on integration and alignment, often overlooking the rich and diverse information embedded in the unique characteristics of each modality.
Figure 2: Overview of MIRROR. WSIs are first partitioned into patches, which are processed through a pre-trained patch encoder to extract patch-level feature representations. These features are subsequently aggregated by the slide encoder to encapsulate slide-level characteristics into a [CLS] token while projecting patch embeddings into the shared pathological latent space. Transcriptomics data are preprocessed using RFE and manual selection to identify high disease-related genes. The refined transcriptomic features are then embedded into a compact representation and mapped into the shared latent space via an RNA encoder. An alignment module for each modality aligns representations across modalities, guided by the alignment loss ($L_{align}$). Meanwhile, modality-specific retention modules utilize perturbed inputs from both encoded patch and transcriptomics features to capture modality-specific intrinsic structures, contributing to the retention loss ($L_{retention}$). Finally, both slide and transcriptomics representations are processed through a style clustering module to learn and compare their pathological styles against learnable cluster centers, with the clustering loss ($L_{cluster}$) used to align consistent pathological styles within the cluster space.
III Methodology
As illustrated in Figure 2, MIRROR consists of four main components, each described in detail below.
III-A Modality Encoders
MIRROR adopts two Transformer-based encoders to effectively project histopathology and transcriptomics data into the shared pathological latent space.
III-A1 Slide Encoder
WSIs are first partitioned into patches, forming an instance bag that is processed through a pre-trained patch encoder to extract patch-level feature representations, denoted as $\mathbf{P}\in\mathbb{R}^{N\times D_{p}}$, where $N$ is the number of patches in the instance bag and $D_{p}$ is the dimensionality of the patch-level representations. These representations are subsequently fed into the slide encoder $f$ to obtain slide-level representations of patch tokens and a global slide [CLS] token:
$\begin{split}\mathbf{S},\mathbf{S}^{\textit{[CLS]}}=f(\mathbf{P}),\end{split}$ (1)
where $\mathbf{S}\in\mathbb{R}^{N\times D}$ represents the slide-level feature embeddings for the patch tokens, $D$ is the dimensionality of the shared latent space, and $\mathbf{S}^{\textit{[CLS]}}$ is the class token capturing global slide-level information. To achieve effective positional encoding and attention, our approach incorporates a modified version of TransMIL [47], which includes two attention blocks and a Pyramid Position Encoding Generator (PPEG) module.
III-A2 RNA Encoder
Transcriptomics profiles are inherently high-dimensional, encompassing a vast number of genes, many of which exhibit redundancy or limited relevance to oncogenic processes. Thus, effective gene selection is crucial for achieving optimal performance. MIRROR employs a hybrid gene selection strategy that integrates both machine learning-driven feature selection with biologically curated gene filtering to address this challenge. To identify the most discriminative genes, recursive feature elimination (RFE) [48] is utilized to determine a highly performant support set. RFE iteratively trains a predictive model $\psi:\mathbb{R}^{D_{g}}\rightarrow\mathbb{R}$ on the raw transcriptomics input matrix $\mathbf{T}_{\text{raw}}\in\mathbb{R}^{N_{s}\times D_{g}}$, where $N_{s}$ denotes the number of samples and $D_{g}$ is the total number of genes. The importance of each gene is quantified by the squared magnitude of the model’s learned coefficients $\boldsymbol{\beta}=[\beta_{1},\beta_{2},\dots,\beta_{D_{g}}]^{\top}$:
$I_{g}=|\beta_{g}|^{2}.$ (2)
At each iteration, the gene with the lowest importance score is eliminated:
$g^{*}=\arg\min_{g}I_{g},$ (3)
and the model is retrained on the reduced feature set until only $K$ features remain, yielding a compact subset of highly informative genes. Additionally, to further ensure interpretability and biological relevance, manually curated genes associated with specific cancer subtypes are selected based on the COSMIC database [49]. This dual strategy ensures a balance between model performance and biological relevancy as shown in Figure 3.
The selected genes, although significantly reduced, remain numerous for direct encoding. To address this, these genes are passed through an embedding layer to reduce their dimensionality, producing a compact representation denoted as $\hat{\mathbf{T}}\in\mathbb{R}^{D_{t}}$, where $D_{t}$ is the dimensionality of the compact representation. To model intricate gene-gene interactions and extract biologically meaningful transcriptomic representations, MIRROR utilizes a Transformer-based RNA encoder [15]. This encoder incorporates a learnable gene encoding token $\mathbf{G}\in\mathbb{R}^{D_{t}}$, which encapsulates the inherent correlations among gene expression patterns. The final encoded transcriptomic representation is computed as:
$\mathbf{T}=g(\hat{\mathbf{T}},\mathbf{G}),$ (4)
where $\mathbf{T}\in\mathbb{R}^{D}$ represents the encoded transcriptomics features in the shared latent space, and $D$ is the dimensionality, matching the output dimensionality of the slide encoder.
Figure 3: Transcriptomics data distributions in the TCGA-NSCLC dataset. The top panel presents a heatmap visualization of transcriptomics data from two subtypes in TCGA-NSCLC, namely TCGA-LUAD and TCGA-LUSC. The data distribution exhibits substantial variability and clear subtype distinction after preprocessing, providing a robust foundation for representation learning. The bottom bar plot highlights the top 10 most variant genes in the TCGA-NSCLC dataset, identified with reference to the COSMIC database, demonstrating extraordinary biological explainability.
III-B Preliminaries
As discussed in Section I, the output of each encoder can be conceptually decomposed into four parts: (1) disease-relevant, modality-shared; (2) disease-relevant, modality-specific; (3) disease-irrelevant, modality-shared; (4) disease-irrelevant, modality-specific. Hence, for the $i$-th paired sample, the outputs of the slide and RNA encoders can be written as
$\begin{split}\mathbf{S}^{i}&=S_{\mathrm{r},s}^{i}+S_{\mathrm{r},u}^{i}+S_{\mathrm{i},s}^{i}+S_{\mathrm{i},u}^{i},\\ \mathbf{T}^{i}&=T_{\mathrm{r},s}^{i}+T_{\mathrm{r},u}^{i}+T_{\mathrm{i},s}^{i}+T_{\mathrm{i},u}^{i},\end{split}$ (5)
where “r” (relevant) vs. “i” (irrelevant) indicates disease relevance, and “s” (shared) vs. “u” (unique) indicates modality traits. While we speak of disease relevance for clarity, in practice these relevant subspaces can capture any shared pathological characteristics that drive meaningful variation in the data. Additionally, for notational conciseness, $\mathbf{S}$ and $\mathbf{S}^{\textit{[CLS]}}$ are treated equivalently in this context.
III-C Modality Alignment Module
The modality alignment module aims to map each encoder output to a common latent space where modality-shared components are systematically brought closer for paired samples, while pusing away irrelevant or mismatched components. To extract the modality-shared information, the encoded representations $\mathbf{S}^{\textit{[CLS]}}$ and $\mathbf{T}$ are mapped into a shared latent space using a modality alignment module for each modality, defined as:
$\begin{split}\mathbf{S_{\text{align}}}&=f_{\text{align}}(\mathbf{S}^{\textit{[CLS]}}),\\ \mathbf{T_{\text{align}}}&=g_{\text{align}}(\mathbf{T}),\end{split}$ (6)
where $\mathbf{S_{\text{align}}},\mathbf{T_{\text{align}}}\in\mathbb{R}^{D}$ are the aligned representations containing modality-shared information defined as:
$\begin{split}\mathbf{S}_{\text{align}}^{i}&=S_{\mathrm{r},s}^{i}+S_{\mathrm{i},s}^{i},\\ \mathbf{T}_{\text{align}}^{i}&=T_{\mathrm{r},s}^{i}+T_{\mathrm{i},s}^{i}.\end{split}$ (7)
To guide the alignment process, an alignment loss function is employed. Inspired by [50], the loss function is formulated as:
$\begin{split}L_{\text{align}}&=-\frac{1}{2B}\sum_{i=1}^{B}\log\left(\frac{\exp(\tau{\mathbf{S}_{\text{align}}^{i}}^{\top}\mathbf{T}_{\text{align}}^{i})}{\sum_{j=1}^{B}\exp(\tau{\mathbf{S}_{\text{align}}^{i}}^{\top}\mathbf{T}_{\text{align}}^{j})}\right)\\ &\quad-\frac{1}{2B}\sum_{i=1}^{B}\log\left(\frac{\exp(\tau{\mathbf{T}_{\text{align}}^{i}}^{\top}\mathbf{S}_{\text{align}}^{i})}{\sum_{j=1}^{B}\exp(\tau{\mathbf{T}_{\text{align}}^{i}}^{\top}\mathbf{S}_{\text{align}}^{j})}\right),\end{split}$ (8)
where $B$ is the batch size, and $\tau$ is a temperature hyperparameter.
This loss function inherently encourages high similarity between paired samples $(\mathbf{S}_{\text{align}}^{i},\mathbf{T}_{\text{align}}^{i})$ while reducing the similarity between negative pairs $(\mathbf{S}_{\text{align}}^{i},\mathbf{T}_{\text{align}}^{j})$ for $i\neq j$. At convergence, the following conditions hold:
$\mathbf{S}_{\text{align}}^{i}\approx\mathbf{T}_{\text{align}}^{i}\quad\text{and}\quad\mathbf{S}_{\text{align}}^{i}\cdot\mathbf{T}_{\text{align}}^{j}\rightarrow 0\quad\forall i\neq j.$ (9)
where the dot product measures similarity, ensuring that only corresponding pairs exhibit high alignment while unrelated pairs are orthogonal in the shared latent space.
Through this iterative optimization process, the model fosters the formation of distinct clusters, wherein samples sharing similar pathological signatures naturally coalesce in the latent space. Concurrently, the enforced separation between dissimilar instances ensures the emergence of well-differentiated clusters, thereby enhancing the model’s capacity to capture discriminative features.
III-D Modality Retention Module
In contrast to the conventional SSL scenarios where input modalities often exhibit highly overlapping or shared features, histopathology and transcriptomics data are fundamentally heterogeneous. This heterogeneity means that simply aligning these two modalities risks discarding essential modality-specific information encoded in $S_{\mathrm{r},u}^{i},S_{\mathrm{i},u}^{i},T_{\mathrm{r},u}^{i}$, and $T_{\mathrm{i},u}^{i}$. To address this, MIRROR introduces a modality retention module that is explicitly designed to safeguard unique attributes of each modality while allowing the alignment module to focus on shared, cross-modal features.
For histopathology, MIRROR employs a masked patch modeling task. A subset of patch tokens from $\mathbf{S}$ is randomly masked, producing a corrupted representation $\mathbf{S}_{\text{masked}}$, from which the retention module tries to reconstruct the missing tokens by learning the mapping:
$\mathbf{S}_{\text{retention}}=f_{\text{retention}}(\mathbf{S}_{\text{masked}}).$ (10)
The reconstructed representation encapsulates both disease-relevant and irrelevant information:
$\mathbf{S}_{\text{retention}}=S_{\mathrm{r},u}+S_{\mathrm{i},u}.$ (11)
For transcriptomics data, a novel masked transcriptomics modeling task is introduced. A subset of gene representations from $\mathbf{T}$ is randomly masked, forming $\mathbf{T}_{\text{masked}}$. The transcriptomics retention module then reconstructs the masked representations by learning:
$\mathbf{T}_{\text{retention}}=g_{\text{retention}}(\mathbf{T}_{\text{masked}}).$ (12)
Similarly, the retention incorporates both modality-specific and instance-specific information:
$\mathbf{T}_{\text{retention}}=T_{\mathrm{r},u}+T_{\mathrm{i},u}.$ (13)
To ensure the retention of modality-specific information, the representations produced by each encoder should encapsulate rich and holistic semantic information. Consequently, the loss function for modality retention is defined as:
$\begin{split}L_{\text{retention}}&=\frac{1}{2B}\sum_{i=1}^{B}\text{sim}(\mathbf{S}^{i},\mathbf{S}_{\text{retention}}^{i})\\ &+\frac{1}{2B}\sum_{i=1}^{B}\text{sim}(\mathbf{T}^{i},\mathbf{T}_{\text{retention}}^{i}),\end{split}$ (14)
where $\text{sim}(\cdot)$ denotes the similarity measurement function, for which Mean Square Error (MSE) is used. Given the stochastic nature of masking, where any patch or gene can be masked at random, the encoder is compelled to distribute modality-specific intrinsic structures uniformly across the entire representation. This strategy ensures each modality’s unique biological and morphological characteristics remain well-preserved, improving overall fidelity of the learned representations.
TABLE I: Cancer subtype classification on TCGA-BRCA and TCGA-NSCLC.
Dataset Backbone Setting ABMIL PORPOISE Linear Classifier TANGLE MIRROR
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
TCGA- NSCLC ResNet-50 10-shot 0.633 0.572 0.742 0.699 0.660 0.658 0.909 0.909 0.930 0.930
±0.084 ±0.147 ±0.182 ±0.242 ±0.024 ±0.025 ±0.033 ±0.033 ±0.006 ±0.006
All Data 0.870 0.868 0.986 0.986 0.845 0.843 0.989 0.987 0.992 0.992
±0.015 ±0.016 ±0.007 ±0.007 ±0.014 ±0.015 ±0.009 ±0.009 ±0.011 ±0.011
Phikon 10-shot 0.709 0.703 0.786 0.767 0.699 0.694 0.925 0.924 0.939 0.938
±0.030 ±0.029 ±0.155 ±0.183 ±0.030 ±0.029 ±0.036 ±0.039 ±0.017 ±0.017
All Data 0.906 0.905 0.984 0.992 0.901 0.900 0.982 0.982 0.994 0.994
±0.010 ±0.010 ±0.007 ±0.008 ±0.008 ±0.008 ±0.012 ±0.012 ±0.007 ±0.007
TCGA- BRCA ResNet-50 10-shot 0.796 0.480 0.902 0.726 0.862 0.507 0.910 0.757 0.910 0.786
±0.075 ±0.028 ±0.010 ±0.152 ±0.031 ±0.076 ±0.036 ±0.121 ±0.032 ±0.092
All Data 0.860 0.462 0.923 0.894 0.901 0.762 0.956 0.909 0.958 0.902
±0.030 ±0.001 ±0.011 ±0.022 ±0.029 ±0.020 ±0.018 ±0.029 ±0.017 ±0.039
Phikon 10-shot 0.712 0.589 0.902 0.731 0.758 0.613 0.921 0.829 0.923 0.832
±0.131 ±0.078 ±0.030 ±0.155 ±0.059 ±0.065 ±0.031 ±0.068 ±0.025 ±0.051
All Data 0.919 0.825 0.934 0.899 0.924 0.821 0.952 0.887 0.955 0.902
±0.008 ±0.021 ±0.030 ±0.009 ±0.038 ±0.092 ±0.022 ±0.063 ±0.023 ±0.047
TABLE II: Cancer survival analysis on TCGA-BRCA and TCGA-NSCLC datasets.
Dataset Backbone Setting ABMIL PORPOISE Linear Classifier TANGLE MIRROR
[…]
III-E Style Clustering Module
To mitigate redundancy and enhance the representation of disease-relevant information, MIRROR incorporates a style clustering module designed to project both $\mathbf{S}^{\textit{[CLS]}}$ and $\mathbf{T}$ into a shared latent space. Within this module, the latent representations for each sample, denoted as $z_{S}$ and $z_{T}$, are regularized to follow a standard normal distribution $\mathcal{N}(0,I)$. This regularization is enforced by minimizing the Kullback–Leibler (KL) divergence [51]:
$L_{\text{style}}\;=\;\mathrm{KL}\!\bigl{(}z_{S}\,\big{\|}\,\mathcal{N}(0,I)\bigr{)}\;+\;\mathrm{KL}\!\bigl{(}z_{T}\,\big{\|}\,\mathcal{N}(0,I)\bigr{)},$ (15)
where the minimization of $L_{\text{style}}$ suppresses unnecessary degrees of freedom in the latent representation, thereby reducing redundancy and promoting a more compact encoding.
To facilitate the alignment of disease-relevant features across modalities, MIRROR introduces a set of learnable cluster centers, $\mathbf{P}\in\mathbb{R}^{D}$, which are normalized to prevent degeneracy. Each latent representation $z_{S}$ and $z_{T}$ is softly assigned to these centers, producing assignment distributions $\mathbf{S}_{\text{cluster}}$ and $\mathbf{T}_{\text{cluster}}$, respectively. The alignment between these assignments is encouraged through a bidirectional KL divergence penalty:
$L_{\mathrm{cluster}}\;=\;\mathrm{KL}\!\bigl{(}\mathbf{S}_{\text{cluster}}\,\|\,\mathbf{T}_{\text{cluster}}\bigr{)}\;+\;\mathrm{KL}\!\bigl{(}\mathbf{T}_{\text{cluster}}\,\|\,\mathbf{S}_{\text{cluster}}\bigr{)},$ (16)
ensuring mutual consistency between the two modalities.
While both $z_{S}$ and $z_{T}$ are constrained to follow the standard normal distribution $\mathcal{N}(0,I)$, only the disease-relevant consistent pathological signatures of these representations are expected to exhibit meaningful correlations across paired samples. Specifically, pathologically significant phenotypes derived from histopathology data and cancer-related gene expressions or pathways in transcriptomics data are expected to exhibit consistent patterns across modalities. Conversely, variations that are not related to the disease, such as differences in tissue structure that do not correlate with pathological conditions or fluctuations in the expression of non-cancer-related genes, are inherently variable and lack consistency. As a result, these irrelevant factors remain unaligned and do not contribute to minimizing the loss function.
Formally, let $z_{S}$ and $z_{T}$ be $d$-dimensional latent vectors such that:
$z_{S}=\bigl{[}z_{S}^{\mathrm{r}},\;z_{S}^{\mathrm{i}}\bigr{]},\quad z_{T}=\bigl{[}z_{T}^{\mathrm{r}},\;z_{T}^{\mathrm{i}}\bigr{]},$ (17)
where $z_{S}^{\mathrm{r}}$ and $z_{T}^{\mathrm{r}}$ encode disease-relevant features, while $z_{S}^{\mathrm{i}}$ and $z_{T}^{\mathrm{i}}$ capture modality-specific or irrelevant variations. The joint minimization of the style and clustering objectives:
$L_{\text{style clustering}}=L_{\text{style}}+L_{\text{cluster}},$ (18)
encourages the alignment of $z_{S}^{\mathrm{r}}$ and $z_{T}^{\mathrm{r}}$ by fostering similar cluster assignments while suppressing correlations between the irrelevant components $z_{S}^{\mathrm{i}}$ and $z_{T}^{\mathrm{i}}$.
In summary, the style clustering module effectively disentangles disease-relevant information from irrelevant variability, ensuring that only meaningful features are learned across modalities.
III-F Global Optimization Objective
The global optimization objective is formally defined as:
$L=L_{\text{align}}+L_{\text{retention}}+L_{\text{style clustering}}.$ (19)
Minimizing this composite loss function promotes the enrichment of disease-relevant information, capturing both modality-shared and modality-specific components, denoted by $S_{r,s}$, $S_{r,u}$, $T_{r,s}$, and $T_{r,u}$. Simultaneously, it suppresses disease-irrelevant information, represented by $S_{i,s}$, $S_{i,u}$, $T_{i,s}$, and $T_{i,u}$. This selective optimization ensures that the learned feature representations concentrate on patterns essential for accurate disease characterization across modalities, thereby improving the model’s diagnostic and prognostic effectiveness.
Figure 4: Visualization of slide encoder attention weights on TCGA-BRCA and TCGA-NSCLC. Regions exhibiting higher attention scores predominantly correspond to malignant, tumor-bearing tissue, whereas areas with lower scores typically indicate normal regions.
IV Experiments and Results
IV-A Datasets
Two cohorts from the publicly available TCGA dataset [16] were used for evaluation: Breast Invasive Carcinoma (BRCA) and Non-Small Cell Lung Cancer (NSCLC). The TCGA-BRCA cohort comprises two primary subtypes: Invasive Ductal Carcinoma (IDC) with 787 samples and Invasive Lobular Carcinoma (ILC) with 168 samples, totaling 955 samples. Similarly, the TCGA-NSCLC cohort includes two subtypes: Lung Adenocarcinoma (LUAD) with 541 samples and Lung Squamous Cell Carcinoma (LUSC) with 512 samples, totaling 1053 samples. Whole Slide Images (WSIs) from both cohorts were processed into patches at 20×20$\times$20 × magnification, and feature representations were extracted using ResNet-50 [52] and Phikon [53]. Corresponding transcriptomics data were collected from Xena [54] and preprocessed using the proposed novel pipeline, reducing the original 198,620 genes to 40,146 for the BRCA cohort and 10,234 for the NSCLC cohort, forming the novel refined transcriptomics datasets.
IV-B Implementation Details
IV-B1 Data Preprocessing
For WSIs, adhering to the conventional preprocessing protocol established in [55], the foreground tissue regions are first segmented using Otsu’s method [56]. These segmented regions are then divided into patches, forming an instance bag that is processed by a pre-trained patch encoder to extract feature representations. To enable batched training, we perform random sampling to select a fixed number of feature representations, employing replacement if the available number of representations is less than the required count.
For raw transcriptomics data, we first apply RFE with 5-fold cross-validation for each cohort to identify the most performant support set for the subtyping task. To enhance interpretability from a biological perspective, we manually incorporate genes associated with specific cancer subtypes based on the COSMIC database [49], resulting in a one-dimensional transcriptomics feature vector.
IV-B2 Experiment Setup
Our experiments evaluate the efficacy of the proposed model, MIRROR, on cancer subtyping and survival analysis tasks by concatenating the outputs from the two modality-specific encoders and feeding the combined vector into a linear classifier. The evaluations utilize few-shot learning and linear probing with 5-fold cross-validation. To assess generalization ability, two different pre-trained patch encoders are used for evaluation.
The model is optimized using the Adam optimizer [57] with a learning rate of $2\times 10^{-5}$. Implementation is conducted in PyTorch [58], and the model is trained for 100 epochs with a batch size of $16$. All training sessions are performed on a single NVIDIA GeForce RTX 3090 GPU paired with an Intel i7-12700K CPU and 32GB of memory.
IV-C Downstream Tasks
Two tasks are used for downstream evaluations, namely subtype classification and survival analysis. We compared the proposed MIRROR against ABMIL [59], PORPOISE [60], Linear Classifier, and TANGLE [11], utilizing two distinct backbones under both 10-shot and linear probing settings. Among these methods, ABMIL exclusively takes histopathology as input, whereas the other models integrate both histopathology and transcriptomics data. To ensure fair comparisons and optimal performance, all methods are trained on the proposed transcriptomic datasets. Notably, the original TANGLE framework evaluates only histopathology features. To comprehensively assess its performance and maintain methodological consistency, we extend TANGLE by incorporating both histopathology and transcriptomic features, aligning its input with that of MIRROR.
IV-C1 Subtype Classification
Accuracy and F1 score are employed as performance metrics to evaluate the classification effectiveness of the compared methods.
As illustrated in Table I, the proposed MIRROR consistently achieves superior performance, demonstrating its ability to effectively integrate histopathology and transcriptomic data for enhanced classification accuracy. On the TCGA-NSCLC dataset, MIRROR achieves the highest accuracy and F1 score across all settings, outperforming all baseline methods. Notably, its strong performance remains consistent regardless of the chosen backbone architecture, highlighting its robustness on the TCGA-NSCLC dataset. Similarly, on the TCGA-BRCA dataet, MIRROR achieves state-of-the-art results, surpassing the second best model’s (TANGLE) result by 1.5% in F1 score under linear probing setting with the Phikon backend. These findings validate the robustness and generalizability of MIRROR across different cancer subtypes and backbone architectures.
IV-C2 Survival Prediction
To assess the performance of our model in survival analysis, we employ a discrete survival model that categorizes patients into four distinct bins. The Concordance Index (C-index) is used as the primary metric to assess performance.
As shown in Table II, the proposed MIRROR achieves significant performance improvements over all baseline methods. On the TCGA-NSCLC dataset, MIRROR surpasses PORPOISE by $4.0\%$ and achieves a $7.6\%$ improvement compared to ABMIL under the 10-shot setting with ResNet-50 as the backbone. This enhancement highlights the superior capability of our model to leverage the integrated multimodal features effectively. Additionally, MIRROR consistently outperforms TANGLE across all settings, further emphasizing its robust feature representation and generalizability. These results highlight the effectiveness of MIRROR in tackling complex survival prediction tasks, reinforcing its potential for improving prognostic modeling in computational pathology.
IV-D Ablation Study
TABLE III: Ablation Study of Model Components on TCGA-BRCA.
Module Setting Performance
Alignment Retention Clustering Subtyping Survival
0.953 $\pm$ 0.024 0.612 $\pm$ 0.075
0.954 $\pm$ 0.028 0.610 $\pm$ 0.085
0.950 $\pm$ 0.016 0.654 $\pm$ 0.083
0.955 $\pm$ 0.023 0.665 $\pm$ 0.082
We conducted an ablation study to evaluate the effectiveness of the key components of MIRROR, including the modality alignment module, modality retention module, and style clustering module. The study was performed on the TCGA-BRCA dataset using linear probing with Phikon as the backbone, comparing MIRROR’s performance both with and without these components.
As shown in Table III, the results highlight the contributions of each module. The modality retention module, while leading to a slight decline in survival analysis performance, enhances subtype classification, demonstrating its ability to preserve modality-specific information. Additionally, the introduction of the style clustering module encourages the model to focus on disease-relevant patterns while mitigating redundancy, enabling the extraction of more informative features. This refinement is reflected in the improved feature retention performance observed during training. With all modules incorporated, MIRROR achieves the best overall performance, confirming the synergistic benefits of the proposed components.
IV-E Visualization
Figure 5: Visualization of transcriptomics features encoded by MIRROR on the TCGA-NSCLC dataset, compared to those produced by TANGLE. Orange dots represent samples from TCGA-LUSC, while blue dots correspond to TCGA-LUAD. MIRROR clearly yields more distinct and representative feature distributions.
Here, we demonstrate the explainability of MIRROR with attention weights visualization from the slide encoder and UMAP [61] projections of features extracted by RNA encoder.
As illustrated in Figure 4, we selected two representative samples from TCGA-BRCA and TCGA-NSCLC, respectively. The attention weights from the slide encoder are visualized using heatmaps to highlight regions of interest within the images. Areas shown in warmer colors indicate regions which the model pays more attention to. The results suggest MIRROR consistently attends to disease-relevant regions, particularly on regions identified as cancerous. This strong alignment showcases MIRROR’s extraordinary interpretability, confirming that it not only excels in downstream tasks but also provides transparent, biologically grounded decision-making insights.
Additionally, as depicted in Figure 5, we visualize features extracted by RNA encoders from TANGLE and MIRROR on the TCGA-NSCLC dataset. Orange and blue dots represent samples from TCGA-LUSC and TCGA-LUAD, respectively. Features generated by MIRROR exhibit significantly improved class separability, with minimal overlap between TCGA-LUSC and TCGA-LUAD clusters. In contrast, TANGLE’s feature space demonstrates considerable inter-class mixing, suggesting weaker discriminative capability. These findings highlight the advantage of MIRROR in generating more representative and structured feature distributions, ultimately contributing to superior model performance on downstream tasks.
V Conclusion
In this paper, we introduced MIRROR, a novel multi-modal pathological SSL framework designed for joint pre-training of histopathology and transcriptomics data. Our approach is tailored to align modality-shared information while retaining modality-specific unique features. Additionally, it employs a style clustering module to reduce redundancy and preserve disease-relevant representations. Furthermore, we proposed a novel transcriptomics data preprocessing pipeline to efficiently identify disease-related genes, resulting in refined, disease-focused transcriptomics datasets. The proposed method was trained and evaluated on two distinct cohorts from the TCGA dataset, demonstrating superior performance and underscoring the efficacy of our design. Future work includes expanding evaluations to additional tasks and cohorts, as well as developing more advanced approaches for extracting information from raw transcriptomics data.